OpenStudy (anonymous):

solve for k, show work 5/7 ( k + 5 ) = -7

OpenStudy (kamille):

5/7(k+5)=-7|*7/5 k+5=-49/5 k+5=-9.8|-5 k=-14.8

OpenStudy (anonymous):

k + 5 = 9.8| -5?